### v2

@@ 5.3.2 Perception and Available Resources Across Different Academic Focuses
We have also examined how different academic focuses of the universities affect their perceptions and resource provision regarding GenAI tools. We divide the top 100 universities into two groups. Institutions recognized for their programs in technology, engineering, and science, such as the California Institute of Technology, are categorized as “technology-oriented universities”. Conversely, universities known for their broader and comprehensive academic subjects, covering areas such as arts, social sciences, and humanities alongside other fields are classified as “comprehensive universities”, such as Harvard University.
An evident trend emerges when comparing the perceptions and resources across the two groups (see Fig. 11). The technology-oriented universities show a higher level of caution and careful consideration in their policies along with more comprehensive and diverse guidelines and resources, which reflect an active engagement with GenAI’s intricacies and potential complex implications in their academic contexts. For example, programming homework can be prevalent in some STEM courses, so there might be more concerns about academic dishonesty due to AI’s ability to provide definitive solutions [21]. On the other hand, there are a number of comprehensive universities that tend to adopt a more welcome, supportive, and positive stance regarding the use of ChatGPT. This variation may indicate that the academic specialization of a university might affect its approach to managing and integrating GenAI in its academic education settings and also suggest the necessity for universities to develop more discipline-specific policies and resources that are sensitive to the unique contexts and challenges of different academic domains in the future.
@@ 6 Discussion
The integration of ChatGPT and other GenAI tools into educational contexts has been mixed with enthusiasm and concerns. The analysis of policies, statements, guidelines and resources regarding ChatGPT of the top 100 universities in the US shows that most of the universities tend to approach the situation with careful consideration and a rich diversity of teaching support, which reflects the increased awareness and efforts from higher education institutions. Some were in a period of observation, evaluating more impacts of ChatGPT on the educational contexts or waiting for the approval of updated university policies [35], which may indicate the potential emergence of clearer policies and guidance in the future.
@@ 6.1 Recommendations for Educators to Make Policies and Guidelines
For teacher trainers and educators who make policies and provide guidelines, this study suggests that it is crucial to establish clear policies and guidelines with the consideration of discipline-specific contexts. While many universities allow instructors to decide their own course policies regarding the use of GenAI, it can be challenging for instructors as they have to not only understand the needs of their students but also align that with the ongoing changes of these tools [5, 40]. To help with establishing policies and guidelines, policymakers can engage with educators across various departments to understand how they currently teach and prepare to teach, what assignments and activities their students are working on, and areas they need to be aware of. Developing policies and resources should be a shared effort that is designed with consideration of a range of academic domains. Depending on some programs’ special contexts and needs, teacher trainers and educators should be encouraged to discuss, evaluate, shape, and refine the policies to ensure the rules align with and support the specific needs of their teaching contexts.
@@
The results also recommend taking precautions when managing sensitive or proprietary information, whether faculty their own or that of their students. Teaching resources and teacher trainers can prepare more training workshops to explicitly discuss the inherent privacy risks of GenAI and how to protect student’s personal information if ChatGPT is integrated into teaching, especially in some activities such as grading and providing feedback. Guidelines and resources should clearly outline what types of information that are safe to share with the AI model and what types are not. On the other hand, academic integrity should extend beyond teaching to include the principles, behaviors, and ethics observed in research as well [20]. More explicit policies and guidelines are necessary to raise researchers’ awareness of which information is considered sensitive and personal and to address the appropriate boundaries for using GenAI in research in higher education.
6.2 Implications for Educators in Teaching Practices
For educators in teaching practice, the study offers a comprehensive overview of the primary resources regarding using GenAI in higher education. Despite the potential risks posed by GenAI, we find that numerous universities have introduced the potential benefits of GenAI and proposed pedagogical applications that enable instructors to leverage ChatGPT effectively in their teaching preparation and practice. With appropriate guidance, GenAI can serve as a helpful and powerful tool for teachers in activity development, implementation, and assessment [12, 23, 26]. Given the practical impossibility of prohibiting ChatGPT use among university students [35], teachers and educators have come to accept, adapt, and embrace its presence [24] and actively engage with university resource centers, seeking new techniques and approaches to incorporate GenAI to enhance student learning.
As a majority of universities give instructors agency to regulate and incorporate the use of GenAI in their own classes, it is important for instructors to decide and apply aligning with their subjects’ specific contexts and learning objectives. The finding offers major pedagogical strategies that can be insightful for teaching with ChatGPT, such as using AI tools to generate writing prompts, grading rubrics, or quiz questions based on the course materials, asking students to compare/contrast AI-generated texts with human writing, and asking students to revise and edit AI-generated information. To incorporate GenAI into classrooms, we suggest instructors explore the pros and cons of GenAI and consider how it can enhance or detract from their subject-specific teaching environments. It is crucial to reflect on the specific student learning outcomes and how GenAI can be used to achieve these goals. For example, in a genre-based writing class, ChatGPT might be employed as a tool for providing writing samples for different contexts, purposes, and audiences, thereby demonstrating different writing conventions and tones. Conversely, in a course aimed at developing research skills, instructors might advise against dependence on ChatGPT for providing article sources, emphasizing instead the importance of students’ critical evaluation of sources and original thought.
To address the concerns about plagiarism and academic dishonesty in teaching and learning, teachers should consider updating their curriculum and evolving activities to avoid students’ misuse of GenAI. First, we suggest instructors establish clear policies and guidelines to explicitly specify students with what they can and cannot do with GenAI in the course syllabi. Given that numerous universities have already offered templates or examples for syllabus language, developing the policies and guidelines should be straightforward. Second, having in-class discussion activities about the ethical use of GenAI and its impact on academic integrity can also help students understand the difference between collaboration with AI and plagiarism. Third, instructors can consider evolving their curriculum and activities by applying some of the prevention strategies discussed in the findings, such as asking students to connect with their personal knowledge and experiences, provide proper academic citations, and reference class materials that are unavailable online. It would also be helpful for instructors to divide their assignments into smaller steps (brainstorming, first draft, second draft, peer review, reflection) to allow students to engage in learning throughout the course instead of relying on one big submission at the end of their course. The discussion and activities can not only prepare students to navigate the complexities of technology but also promote a deeper understanding of the responsibilities that come with using such tools.
Last but not least, considering the current AI detection tools are not reliable and supported to determine plagiarism by most universities [8, 38], teachers should adopt multifaceted evaluation strategies to assess students’ work. One practical approach can be to assess students’ written work based on their previous and in-class performance. AI-generated content cannot fully replicate the unique writing style of individual students, which often includes word selection, phrase usage, language patterns, and students’ personal insights developed over their learning. Comparing multiple works for consistency may be a way to identify AI-generated content. Some other prevention strategies, such as asking students to explain their answers in multimodal ways, such as presentations and audio threads, discussed in the results section can also be applied to avoid the inappropriate use of AI. It is important to note that instructors should treat each student’s work as a learning opportunity, thus, providing feedback at multiple levels and asking students to reflect on how they make changes to their assignments can be beneficial in both areas (preventing the excessive use of AI and students’ gaining knowledge). In addition, accusing students of using AI needs to be addressed with caution and care, in order to maintain a trusting relationship between students and their instructors. More guidance on how to establish these harder conversations between faculty and students needs to be discussed and developed further.
@@ 7 Conclusion
This study delves into the academic policies, resources, and guidelines of the top 100 U.S. universities regarding ChatGPT and other GenAI tools in higher education. Data was collected from publicly available official university sources, such as the Office of Provost and the Center of Teaching and Learning. The results reveal a prevalent balanced and open yet cautious and thoughtful attitude toward integrating AI technology given concerns mainly on ethical issues, inherent limitations, and data privacy. A number of universities encourage instructors to develop their own policies and guidelines for the use of ChatGPT, respecting the specific contexts and needs of their disciplines. For resources and guidelines, the most popular teaching support includes syllabus samples and templates, workshops, articles, and individual consultations with topics on technical introduction, pedagogical applications, prevention strategies, limitations, and detection tools, to help instructors adapt their teaching practices in the age of GenAI. These efforts highlight a variety of opportunities and challenges presented by GenAI along with raising the necessity to further enrich and refine the current guidelines as well as curriculum for instructors, students, and researchers in higher education.
In the era of GenAI, what we should not do is do nothing. Actively engaging with these technology advancements is significant for better leveraging their potential and effectively mitigating their risks. The findings of this study may have important implications for educators in the contexts of both policy-making and teaching practices. For educators who need to make policies for their own classes and/or departments, we recommend designing policies according to their discipline-specific contexts and taking precautions when managing sensitive information. For educators in teaching practices, the pedagogical implications include accepting, adapting, and embracing the presence of GenAI, aligning its use with specific learning objectives, updating curriculum to guide and prevent students from misuse, as well as applying multifaceted evaluation strategies instead of relying on AI detectors.
[…]
Data availability
The authors declare that the data supporting the findings of this study are available within the paper.
@@
[…]
Confict of interest The authors have no conficts of interest to declare that are relevant to the content of this article.

### v3

@@ 5.3.2 Perception and Available Resources Across Different Academic Focuses
We have also examined how different academic focuses of the universities affect their perceptions and resource provision regarding GenAI tools. We divide the top 100 universities (n = 104) into two groups. Institutions recognized for their programs in technology, engineering, and science, such as the California Institute of Technology, are categorized as “technology-oriented universities” (n = 24). Conversely, universities known for their broader and comprehensive academic subjects, covering areas such as arts, social sciences, and humanities alongside other fields are classified as “comprehensive universities” (n = 80), such as Harvard University.
An evident trend emerges when comparing the perceptions and resources across the two groups (see Fig. 11). The technology-oriented universities show a higher level of caution and careful consideration in their policies along with more comprehensive and diverse guidelines and resources, which reflect an active engagement with GenAI’s intricacies and potential complex implications in their academic contexts. For example, programming homework can be prevalent in some STEM courses, so there might be more concerns about academic dishonesty due to GenAI’s ability to provide definitive solutions (Michel-Villarreal et al., 2023). On the other hand, there are a number of comprehensive universities that tend to adopt a more welcome, supportive, and positive stance regarding the use of GenAI. This variation may indicate that the academic specialization of a university might affect its approach to managing and integrating GenAI in its academic education settings. Considering the nature of different academic disciplines and contexts, the results highlight the importance of developing discipline-specific policies and guidelines that can address challenges and accommodate the needs of various academic domains. It can also reflect that the integration of GenAI in educational practice needs to align with the objectives of each discipline to effectively enhance student learning experiences.
@@ 6 Discussion
The integration of ChatGPT and other GenAI tools into educational contexts has been mixed with enthusiasm and concerns. In order to incorporate GenAI tools in higher education in a more ethical and responsible way, this study analyzes the currently available policies, statements, guidelines and resources regarding GenAI, especially ChatGPT, of the top 100 universities in the U.S.. The results show that most of the universities tended to approach the situation with careful consideration and a rich diversity of teaching support, which reflects the increased awareness and efforts from higher education institutions. Some were in a period of observation, evaluating more impacts of GenAI on the educational contexts or waiting for the approval of updated university policies (Sullivan et al., 2023).
6.1 Implications for Educators in Teaching Practices
For educators in teaching practice, the study offers a comprehensive overview of the primary resources regarding using GenAI in higher education. Despite the potential risks posed by GenAI, we find that numerous universities have introduced the potential benefits of GenAI and proposed pedagogical applications that enable instructors to leverage ChatGPT effectively in their teaching preparation and practice. With appropriate guidance, GenAI can serve as a helpful and powerful tool for teachers in activity development, implementation, and assessment (Hodges and Kirschner, 2024; Mishra et al., 2023; Oravec, 2023). Given the practical impossibility of prohibiting ChatGPT use among university students (Sullivan et al., 2023), teachers and educators have come to accept, adapt, and embrace its presence (Moorhouse et al., 2023) and actively engage with university resource centers, seeking new techniques and approaches to incorporate GenAI to enhance student learning.
As a majority of universities give instructors agency to regulate and incorporate the use of GenAI in their own classes, it is important for instructors to decide and apply to align with their subjects’ specific contexts and learning objectives. The finding offers major pedagogical strategies that can be insightful for teaching with ChatGPT, such as using GenAI tools to generate writing prompts, grading rubrics, or quiz questions based on the course materials, asking students to compare/contrast AI-generated texts with human writing, and asking students to revise and edit AI-generated information. To incorporate GenAI into classrooms, we suggest instructors explore the pros and cons of GenAI and consider how it can enhance or detract from their subject-specific teaching environments. It is crucial to reflect on the specific student learning outcomes and how GenAI can be used to achieve these goals. For example, in a genre-based writing class, ChatGPT might be employed as a tool for providing writing samples for different contexts, purposes, and audiences, thereby demonstrating different writing conventions and tones. Conversely, in a course aimed at developing research skills, instructors might advise against dependence on ChatGPT for providing article sources, emphasizing instead the importance of students’ critical evaluation of sources and original thought.
To address the concerns about plagiarism and academic dishonesty in teaching and learning, teachers should consider updating their curriculum and evolving activities to avoid students’ misuse of GenAI. First, we suggest instructors establish clear policies and guidelines to explicitly specify to students what they can and cannot do with GenAI in the course syllabi. Given that numerous universities have already offered templates or examples for syllabus language, developing the policies and guidelines should be straightforward. Second, having in-class discussion activities about the ethical use of GenAI and its impact on academic integrity can also help students understand the difference between collaboration with GenAI and plagiarism. Third, instructors can consider evolving their curriculum and activities by applying some of the prevention strategies discussed in the findings, such as asking students to connect with their personal knowledge and experiences, providing proper academic citations, and referencing class materials that are unavailable online. It would also be helpful for instructors to divide their assignments into smaller steps (brainstorming, first draft, second draft, peer review, reflection) to allow students to engage in learning throughout the course instead of relying on one big submission at the end of their course. The discussion and activities can not only prepare students to navigate the complexities of technology but also promote a deeper understanding of the responsibilities that come with using such tools.
Last but not least, considering the current GenAI detection tools are not reliable and supported to determine plagiarism by most universities (Elkhatat et al., 2023; Weber-Wulff et al., 2023), teachers should adopt multifaceted evaluation strategies to assess students’ work. One practical approach can be to assess students’ written work based on their previous and in-class performance. AI-generated content cannot fully replicate the unique writing style of individual students, which often includes word selection, phrase usage, language patterns, and students’ personal insights developed over their learning. Comparing multiple works for consistency may be a way to identify AI-generated content. Some other prevention strategies, such as asking students to explain their answers in multimodal ways, such as presentations and audio threads, discussed in the results section can also be applied to avoid the inappropriate use of GenAI. It is important to note that instructors should treat each student’s work as a learning opportunity, thus, providing feedback at multiple levels and asking students to reflect on how they make changes to their assignments can be beneficial in both areas (preventing the excessive use of GenAI and students’ gaining knowledge). In addition, accusing students of using GenAI needs to be addressed with caution and care, in order to maintain a trusting relationship between students and their instructors. More guidance on how to establish these harder conversations between faculty and students needs to be discussed and developed further.
@@ 6.2 Recommendations for Educators to Make Policies and Guidelines
For teacher trainers and educators who make policies and provide guidelines, this study suggests that it is crucial to establish clear policies and guidelines with the consideration of discipline-specific contexts. While many universities allow instructors to decide their own course policies regarding the use of GenAI, it can be challenging for instructors as they have to not only understand the needs of their students but also align that with the ongoing changes of these tools (Chiu, 2023; Zastudil et al., 2023). To help with establishing policies and guidelines, policymakers can engage with educators across various departments to understand how they currently teach and prepare to teach, what assignments and activities their students are working on, and areas they need to be aware of. Developing policies and resources should be a shared effort that is designed with consideration of a range of academic domains. Depending on some programs’ special contexts and needs, teacher trainers and educators should be encouraged to discuss, evaluate, shape, and refine the policies to ensure the rules align with and support the specific needs of their teaching contexts.
@@
The results also recommend taking precautions when managing sensitive or proprietary information, whether faculty their own or that of their students. Teaching resources and teacher trainers can prepare more training workshops to explicitly discuss the inherent privacy risks of GenAI and how to protect students’ personal information if ChatGPT is integrated into teaching, especially in some activities such as grading and providing feedback. Guidelines and resources should clearly outline what types of information that are safe to share with GenAI and what types are not. On the other hand, academic integrity should extend beyond teaching to include the principles, behaviors, and ethics observed in research as well (Macfarlane et al., 2014). More explicit policies and guidelines are necessary to raise researchers’ awareness of which information is considered sensitive and personal and to address the appropriate boundaries for using GenAI in research in higher education.
@@ 7 Conclusion
This study delves into the academic policies, resources, and guidelines of the top 100 U.S. universities (n = 104) regarding ChatGPT and other GenAI tools in higher education, and thereby informs educators in teaching practices as well as future policy-making. Data was collected from publicly available official university sources, such as the Office of Provost and the Center of Teaching and Learning. The results reveal a prevalent balanced and open yet cautious and thoughtful attitude toward integrating GenAI technology given concerns mainly on ethical issues, inherent limitations, and data privacy. A number of universities encourage instructors to develop their own policies and guidelines for the use of GenAI (n = 57, 54.8%), respecting the specific contexts and needs of their disciplines. For resources and guidelines, most popular teaching support includes syllabus samples and templates (n = 65, 62.5%), workshops (n = 40, 38.5%), articles (n = 74, 71.1%), and individual consultations with topics on technical introduction (n = 41, 39.4%), pedagogical applications (n = 56, 53.8%), prevention strategies (n = 57, 54.8%), limitations (n = 56, 53.8%), and detection tools (n = 59, 56.7%), to help instructors adapt their teaching practices in the age of GenAI. These efforts highlight a variety of opportunities and challenges presented by GenAI along with raising the necessity to further enrich and refine the current guidelines as well as curriculum in higher education.
In the era of GenAI, what we should not do is to stay stagnant. Actively engaging with these technology advancements is significant for better leveraging their potential and effectively mitigating their risks. The findings of this study may have important implications for educators in the contexts of both teaching practices and policy design. For educators in teaching practices, the pedagogical implications include accepting, adapting, and embracing the presence of GenAI, aligning its use with specific learning objectives, updating the curriculum to guide and prevent students from misuse, as well as applying multifaceted evaluation strategies instead of relying on GenAI detectors. For educators who need to make policies for their own classes and/or departments, we recommend designing policies according to their discipline-specific contexts and take precautions when managing sensitive information.
@@
Declaration of competing interest: The authors declare no financial or personal relationships that could inappropriately influence their work.
Data availability: The authors declare that the data supporting the findings of this study are available within the paper.
Funding: This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Acknowledgements: Not applicable.